OpenStudy (anonymous):

Solve the following system equations: 2x + 4y – 3z = –7 3x + y + 4z = –12 x + 3y + 4z = 4 POSIBBLE ANSWERS: (–6, 2, 1) (6, 2, 1) (6, –2, –1) (–6, –2, –1)

OpenStudy (amistre64):

when in doubt, plug in the answers to see what fits

OpenStudy (anonymous):

there are 2 ways to go about this. you can do trial and error with the 4 different possible solutions or you can actually solve the equations

OpenStudy (amistre64):

trial and error is the easiest way to go lol

OpenStudy (amistre64):

a can tell right now that the last 2 options are off the table

OpenStudy (akshay_budhkar):

lol solving three equations would take your breath away, rather u will solve 4 other questions with trial and error

OpenStudy (amistre64):

the 2nd one falls flat as well, id test the first to be sure

OpenStudy (anonymous):

to solve the equations you need to solve for a variable via 2 equations and then substitute it in the other one. the easiest way to do this is take bottom 2 equations. multiply 2nd one by (-1)

OpenStudy (anonymous):

-6,2,1??? AM i correct?

OpenStudy (anonymous):

the first one satisfy the equation

OpenStudy (amistre64):

dunno, but i think its your best shot

OpenStudy (anonymous):

but u can solve it bu cramer or gous rules

OpenStudy (anonymous):

\[(-1)(3x + y + 4z = –12) = -3x-y-4z=12 add this to bottom eqn to cancel out z's

OpenStudy (amistre64):

taking test and doing math are 2 seperate entities ;)

OpenStudy (mathmate):

For the trial and error, you can add up the equations to make a new one, 6x+8y=5z=-15 and you need to test only once for each choice.

OpenStudy (anonymous):

thank you everyone.. medals for all my helpers